OpenStudy (anonymous):

A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

please help

OpenStudy (campbell_st):

opps should read... multiply every term in the 1st equation by -3.... what is the result

OpenStudy (anonymous):

honestly i have no idea ... idk what your talking about can you explain it to me

OpenStudy (campbell_st):

so its \[-3 \times -3x + -3 \times 7y = -3 \times -16\] can you simplify that

OpenStudy (anonymous):

no.. ): im sorry im stupid and dont understand math one bit

OpenStudy (campbell_st):

oops should be 9x - 21y = 48 does that make sense?

OpenStudy (anonymous):

i guess is that A?

OpenStudy (campbell_st):

yes, that's the equivalent equation.... you did that by multiplying every term by -3... the reason I chose -3... is that you get 9x -21y = 48 when you add the 2 equations x will be eliminated + -9x + 5y = 16 which will allow you to solve for y ---------------- -16y = 64 now solve for y. to get an x value, substitute your answer from y into either equation to get x. hope that helps

OpenStudy (anonymous):

okay i kinda understand.. is that still part a ? or is that to help with b?

OpenStudy (campbell_st):

so the equivalent equation is 9x -21y = 48 then you need to solve the 2 equations, so you can get the values of x and y so that is partly done 1. find y if -16y = 64 2. substitute that value into either of the original equations... to get x Part B use your solutions and substitute them into the equivalent equation. to show the solution is correct for the equivalent equation... hope that makes sense.

im sorry i dont get it )))):